• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 03:26:43 on November 01, 2008

    Entry number 245777

    Followups:

    keyword=target work

    After Yi lowered the EPR RF coil in the oven, Jin/Chiranjib and I continued the work on checking transverse line optics. They look reasonable. The spectrum still shows no absorption dip. Checked the optical fiber for the spectra-analyzer, seems it is easy to rotating around and also it got hot (close to oven). We tried vertical lasers and they seem to have good absorption spectra. So we decided to leave it pumping in vertical direction for now. We used 4 lasers for vertical lines and one laser kept with transverse line (N1-C1). After clean all the optical elements (fiber tips, mirrors etc), we closed the target enclosure and completed the in-hall target work for tonight. Chiranjib adjusted the laser wavelength for the 4 vertical lasers (N7-C3, N8-C4, N9-C5, N10-C6). Per shift worker's phone request (whiteboard instruction) I opened the He3 gas bottle valve before closing the hall for beam. (After I came up to counting house, from whiteboard, seems this could have be done in the next access). Shift worker is instructed to be careful with He3 gas for reference cell run.